OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 1)2 – 16? Identify the x–intercepts by graphing the quadratic equation. i know the y intercept is (0,-15) I'm just not so sure about the x intercept .

OpenStudy (anonymous):

if x=0 y=? if y =0 x=?

OpenStudy (anonymous):

y= (0 -1)^2 -16 y=-15 0 =(x-1)^2 -16 16= (x-1)^2 \[\pm \sqrt{16} =x-1\]

OpenStudy (anonymous):

equate y=0 to get the x equate x=0 to get the y

OpenStudy (anonymous):

finish that up and you'll have two values for x ...the intercepts (roots)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

sure:)